"an aqueous solution containing 10. g of starch per liter has an osmotic pressure of 3.8 mmHg at 25[deg]C. What is the average molar mass of starch?"

I keep getting 48923.68g/mol. We are supposed to use PI=cRT.
1
Expert's answer
2018-10-08T07:52:09-0400

Answer:

PI = 3.8 mmHg = 0,5066 kPa

T=25+273=298 K

R=8.314

Formula:

c= PI / RT

c= 0,5066 / 8.314 x 298

c= 0.000204474381 mol/L

if L=1 to n= 0.000204474381 mol and Mr=m/n

Mr=10/ 0.000204474381=48905.88 g/mol.